# You Only Look at Once for Real-time and Generic Multi-Task

### v1

@@ III-B Decoder
The detection head adopts a decoupled approach that follows the YOLOv8 detect head, using convolutional layers to convert high-dimensional features into class predictions and bounding boxes without the objectness branch. Note that the outputs from the detection head differ between training, validation, and prediction modes. In validation and prediction mode, the processing adheres to the right section of Fig 1. Three different resolutions are used as input, and output one tensor that encompasses both class predictions and their respective bounding boxes. In training mode, the detection head cycle applies convolutional layers to each input, outputs three tensors in total. Each tensor contains class predictions and bounding boxes specific to its resolution. The results are subsequently used to compute the loss function.
Our segmentation head is identical across different segmentation tasks. Specifically, it comprises a series of convolutional layers that extract contextual information and one deconvolution layer to restore the resolution to the original image size. Ultimately, we obtain a pixel-level binary mask that matches the size of the original image. 0 represents the background, and 1 represents the object. Algorithm 2 illustrates the processing of the segment head. Furthermore, the segment head’s streamlined architecture, which comprises just 7,940 parameters, significantly bolsters the model’s deployability.
@@
0:  Number of classes $nc$, channels $ch$, Feature map $x$
@@ III-C Loss Function
We employ an end-to-end training approach with a multi-task loss function. Specifically, our loss function comprises two components: detection and segmentation. The formula is shown as follows:
@@
For detection tasks, the loss function is divided into three main components: the classification loss $\mathcal{L}_{\text{class}}$, the object loss $\mathcal{L}_{\text{obj}}$, and the bounding box loss $\mathcal{L}_{\text{box}}$. The classification loss is quantified using the binary cross-entropy loss, $\mathcal{L}_{\text{BCE}}$, which measures the divergence between predicted probabilities and the ground truth. The object loss adopts the distribution focal loss (DFL) [29], denoted as $\mathcal{L}_{\text{DFL}}$. The DFL captures the position of the bounding box via a generalized distribution. This allows the network to rapidly focus on regions close to the annotated position, thereby amplifying their probabilities. The bounding box loss employs the complete IoU (CIoU) loss [30], represented as $\mathcal{L}_{\text{CIoU}}$. The CIoU loss integrates aspects of overlap, distance, and aspect ratio consistency. As a result, it enables the model to more precisely target the object’s shape, size, and orientation. Thus, the detection loss $\mathcal{L}_{\text{dec}}$ can be represented as:
$\mathcal{L}_{\text{det }}=\lambda_{\text{class }}\mathcal{L}_{\text{BCE }}+\lambda_{\text{obj }}\mathcal{L}_{\text{DFL}}+\lambda_{\text{box }}\mathcal{L}_{\text{CIoU }}$ (2)
where $\lambda_{\text{class }}$, $\lambda_{\text{obj }}$, and $\lambda_{\text{box }}$ are corresponding coefficients.
For the segmentation tasks, we utilize an identical loss function. That means $\mathcal{L}_{\text{segda}}$ has the same formula with $\mathcal{L}_{\text{segll}}$. However, they perform backward separately. We will collectively refer to them as $\mathcal{L}_{\text{seg}}$. The formula is shown as follows:
@@
where $\mathcal{L}_{\text{FL }}$ and $\mathcal{L}_{\text{TL }}$ are focal loss [23] and tversky loss [31]. Both of these are widely used loss functions in segmentation tasks. $\lambda_{\text{FL }}$ and $\lambda_{\text{TL }}$ are corresponding coefficients.
Focal loss offers a robust solution for handling imbalanced samples, ensuring that the model doesn’t become overwhelmingly biased towards the predominant and easily learned class. Instead, it places greater emphasis on challenging and underrepresented areas. Tversky loss is an extension of the Dice loss, introducing two additional parameters ($\alpha$ and $\beta$) to assign distinct weights to false positives and false negatives, thereby enhancing its capability to handle imbalanced tasks.
@@ III-D Training Paradigm
Our training approach represents a departure from classical multi-task learning methods prevalent in panoptic autonomous driving tasks. We utilize an end-to-end training mode, performing backward only once for each batch. This means the entire network is optimized collectively, without freezing specific layers or alternating optimization, thus reducing the training time. Algorithm 3 illustrates the training process of a step-by-step.
@@
0:  Target end-to-end network $F$ with parameters $\Theta$; Training dataset $\tau$; Validation dataset $x_{val}$; Threshold for convergence $\phi$; Total loss function $\mathcal{L}=\mathcal{L}_{\text{det }}+\mathcal{L}_{\text{segda }}+\mathcal{L}_{\text{segll }}$
@@
4:        $\hat{y}\leftarrow F(x)$ $\hat{y}$ is prediction
5:        $\mathcal{L}\leftarrow(y,\hat{y})$
6:        $\Theta\leftarrow\arg\min_{\Theta}\mathcal{L}$
7:     end for
8:     $p\leftarrow F(x_{val})$
9:     if $p=p_{\text{epoch}-50}$ then
10:        return  $F(x;\Theta)$ if the model hasn’t any improvements over the last 50 epochs, then training will stop early.
11:     end if
12:  end for
13:  return  $F(x;\Theta)$
@@ IV EXPERIMENTS AND RESULTS
In this section, we will evaluate our model’s performance and inference time on the BDD100K dataset and compare it against classical methods used in multi-task autonomous driving panoptic perception. Additionally, we present extensive ablation studies and provide an analysis of the experimental results.
@@ IV-A1 Dataset
The BDD100K dataset is a prominent resource in the study of multi-task learning in autonomous driving, including 100k samples and multi-task annotations. Beyond sheer volume, the dataset’s significance is underscored by its multifaceted nature, comprising diverse geographies, environmental contexts, and weather conditions. The advantages ensure that models trained on BDD100K achieve robustness and versatility, making it an ideal choice for our research endeavours. The dataset is divided into three parts: a training set with 70k images, a validation set of 10k images, and a testing set encompassing 20k images. Since the labels for the testing set are not public, we evaluated our model on the validation set. Similar to YOLOP, our detection task focuses solely on “vehicle” detection, encompassing categories including car, bus, truck, and train.
@@ IV-A2 Evaluation Metrics
We utilize Recall and mAP50 as the evaluation metrics for object detection tasks. Both metrics are widely recognized and accepted in the detection task. Recall indicates a model’s capability to detect all object instances from the designated classes accurately. mAP50 is derived by taking the mean Average Precision for all classes at an IoU threshold of 0.5. Notably, Average Precision quantifies the area beneath the precision-recall curve. For the segmentation tasks, similar to YOLOP [15], we utilize mIoU to evaluate the drivable area segmentation task. For the lane line segmentation task, we employ both accuracy and IoU as evaluation metrics. However, due to the number of pixel imbalance between the background and foreground in lane line segmentation, we adopt a more meaningful balanced accuracy metric for evaluation. Traditional accuracy can skew results by favouring classes with a larger number of samples. In contrast, balanced accuracy offers a more equitable metric by considering each class’s accuracy. The formula is as follows:
[…]
@@ IV-A3 Experimental Setup and Implementation
We compare our model with several prominent multi-task models. Due to the scarcity of multi-task models, we also include outstanding single-task or two-task models in our comparison. YOLOP and HybridNet represent the state-of-the-art and are open-source multi-task models in panoptic autonomous driving, specifically tailored for the BDD100K dataset. Faster R-CNN and YOLOv8 are exemplary representations of two-stage and one-stage object detection networks, respectively. Both MultiNet and DLT-Net could address multiple panoptic driving perception tasks, and they’ve demonstrated good performance in object detection and drivable area segmentation on the BDD100k dataset. PSPNet [32] excels in semantic segmentation tasks due to its unparalleled capability to aggregate global information. Given the absence of an appropriate multi-task network handling lane line segmentation on the BDD100K dataset, we compare our model against Enet [33], SCNN, and Enet-SAD [34], which are three leading lane detection networks.
@@
We train our model using the SGD optimizer with a learning rate ($lr$) of 0.01, momentum of 0.937, and a weight decay of 0.0005. Initially, our model undergoes a warm-up training for 3 epochs. During this warm-up phase, the momentum of the SGD optimizer is set to 0.8, and the bias lr is 0.1. In our training process, we adopt a linear learning rate annealing strategy. This strategy helps ensure that the model learns rapidly in the early stages of training and converges more stably in the later stages. Additionally, we resize the original image dimensions from 1280 × 720 to 640 × 640. For loss function coefficients, we set $\lambda_{\text{FL}}=24.0$, $\lambda_{\text{TL}}=8.0$, $\lambda_{\text{obj}}=1.5$, $\lambda_{\text{box}}=7.5$, and $\lambda_{\text{class}}=0.5$. Finally, we train with a batch size of 120 on four RTX 4090 for 300 epochs. For evaluation, we set the confidence threshold at 0.001 and the NMS (Non-Maximum Suppression) threshold at 0.6. For prediction, the confidence threshold is set at 0.25, and the NMS threshold is at 0.45. We follow these settings from YOLOP. As a result, the visualization may slightly differ from the quantitative results.
@@ IV-B Experimental results
In this subsection, we train our model end-to-end and compare its performance against other outstanding methods.
@@ IV-B1 Inference Time
One of the primary challenges in deep learning applications is the inference time. Especially in autonomous driving tasks, require deploying models on edge devices, which usually have limited computational resources. Thus, ensuring models are both lightweight and exhibit real-time becomes paramount. In Table I, we reproduce and test the inference times of YOLOP [^1] https://github.com/hustvl/YOLOP, HybridNet [^2] https://github.com/datvuthanh/HybridNets, YOLOv8333https://github.com/ultralytics/ultralytics, and our model, all tested FPS on a GTX 1080 Ti GPU with 1 and 32 batch size. The FPS calculation approach follows HybridNet. Additionally, we provide the number of parameters for each model as one of the evaluation results.
@@
Model
Parameters
FPS (bs=1)
FPS (bs=32)
YOLOP
7.9M
26.0
134.8
HybridNet
12.83M
11.7
26.9
YOLOv8n(det)
3.16M
102
802.9
YOLOv8n(seg)
3.26M
82.55
610.49
A-YOLOM(n)
4.43M
39.9
172.2
A-YOLOM(s)
13.61M
39.7
96.2
The primary difference between A-YOLOM(n) and A-YOLOM(s) is the complexity of the backbone. A-YOLOM(n) is our lightweight backbone network designed with reduced complexity, making it ideal for deployment on edge devices. A-YOLOM(s) has a more complex backbone, offering more powerful performance with increased time overhead, especially in scenarios with bs=32. Compared to other leading multi-task models and single-task models in panoptic autonomous driving, A-YOLOM(n) stands out as more lightweight and offers more efficiency. Specifically, compared to YOLOP, A-YOLOM(n) has significantly fewer parameters and a higher FPS. It achieves a speed-up of x1.53 at bs=1 and x1.28 at bs=32. That indicated that our model is more efficient. Compared to HybridNet, A-YOLOM(n) has a significant advantage in parameters and speed. Despite having more parameters than HybridNet, A-YOLOM(s) is faster. Due to HybridNet doesn’t meet the requirements for real-time performance in both bs=1 and bs=32, we won’t compare its performance with ours in the subsequent section. YOLOv8 is a single-task model, meaning it can only execute one task within a single model. This is the reason we have listed two versions: one as YOLOv8 (det) and the other as YOLOv8 (seg). Although both of them are faster than other models, including ours, they require deploying three separate models on edge devices, which means combining one detection model with two segmentation models into one edge device, amounting to 9.68M parameters. This is x2.18 more than A-YOLOM(n), putting significant pressure on edge devices. Additionally, YOLOv8’s performance across the three tasks is much inferior to ours, which we will discuss in the upcoming section.

### v4

@@ III-B Decoder
The detection head adopts a decoupled approach that follows the YOLOv8 detect head, using convolutional layers to convert high-dimensional features into class predictions and bounding boxes without the objectness branch. This is an anchor-free detector. The structure is shown as Detect module in Fig 2. Note that the outputs from the detection head vary among training, validation, and prediction modes. In validation and prediction mode, three different resolutions are used as input, and output one tensor encompasses both class predictions and their respective bounding boxes. In training mode, the detection head cycle applies convolutional layers to each input, outputs three tensors in total. Each tensor contains class predictions and bounding boxes specific to its resolution. The results are subsequently used to compute the loss function.
Our segmentation head is identical across different segmentation tasks. The structure is shown as Segment module in Fig 2. Specifically, it comprises a series of convolutional layers that extract contextual information and one deconvolution layer to restore the resolution to the original image size. Ultimately, we obtain a pixel-level binary mask that matches the size of the original image. 0 represents the background, and 1 represents the object. Algorithm 2 illustrates the processing of the segment head. Furthermore, the segment head’s streamlined architecture, which comprises just 7,940 parameters, significantly bolsters the model’s deployability.
@@
0:  Number of classes $nc$; channels $ch$; Feature map $x$
@@ III-C Loss Function
We employ an end-to-end training approach with a multi-task loss function. Specifically, our loss function comprises three components: one detection and two segmentation. The formula is shown as follows:
@@
For detection tasks, the loss function is divided into two main branches: the classification branch and the bounding box branch. The classification branch includes binary cross-entropy loss, denoted as $\mathcal{L}_{\text{BCE}}$. The bounding box branch includes distribution focal loss (DFL) [30], denoted as $\mathcal{L}_{\text{DFL}}$ and complete IoU (CIoU) loss [31], represented as $\mathcal{L}_{\text{CIoU}}$. Thus, the detection loss $\mathcal{L}_{\text{dec}}$ can be represented as:
$\mathcal{L}_{\text{det}}=\lambda_{\text{BCE}}\mathcal{L}_{\text{BCE}}+\lambda_{\text{DFL}}\mathcal{L}_{\text{DFL}}+\lambda_{\text{CIoU}}\mathcal{L}_{\text{CIoU}}$ (2)
where $\lambda_{\text{BCE}}$, $\lambda_{\text{DFL}}$, and $\lambda_{\text{CIoU}}$ are corresponding coefficients.
$\mathcal{L}_{\text{BCE}}=-\left[y_{n}\log x_{n}+\left(1-y_{n}\right)\log\left(1-x_{n}\right)\right]$ (3)
where $x_{n}$ is the predicted classification of each object. $y_{n}$ is the ground truth of each object. $\mathcal{L}_{\text{BCE}}$ measures the classification error between the predicted and the ground truth.
$\begin{gathered}\mathcal{L}_{\mathrm{DFL}}\left(\mathcal{S}_{i},\mathcal{S}_{i+1}\right)=-\left(\left(y_{i+1}-y\right)\log\left(\mathcal{S}_{i}\right)+\left(y-y_{i}\right)\log\left(\mathcal{S}_{i+1}\right)\right)\\ S_{i}=\frac{y_{i+1}-y}{y_{i+1}-y_{i}},S_{i+1}=\frac{y_{i}-y}{y_{i}-y_{i+1}}\\ \end{gathered}$ (4)
where $y$ is the ground truth of the bounding box coordinate, which is a decimal. $y_{i+1}$ is ceiling of ground truth $y$. $y_{i}$ is floor of ground truth $y$. $\mathcal{L}_{\text{DFL}}$ measures the displacement between the predicted and ground truth feature locations to make the predicted bounding boxes close to the actual ones.
$\begin{gathered}\mathcal{L}_{\text{CIoU}}=1-CIoU\\ CIoU=\operatorname{IoU}-\frac{\rho^{2}\left(b,b^{gt}\right)}{c^{2}}-\alpha v\\ v=\frac{4}{\pi^{2}}\left(\arctan\frac{w^{gt}}{h^{gt}}-\arctan\frac{w}{h}\right)^{2}\\ \alpha=\frac{v}{(1-IoU)+v}\end{gathered}$ (5)
where $b$ is the central point of the prediction box, $b^{gt}$ is the central point of the ground truth box. $\rho$ is the Euclidean distance between prediction and ground truth points. $c$ is the diagonal length of the smallest enclosing rectangle of the two boxes. Both $v$ and $\alpha$ are the coefficients used to control the ratio. $h$ is the width and height of the prediction box, $w^{gt}$ and $h^{gt}$ are the width and height of the ground truth box. $\mathcal{L}_{\text{CIoU}}$ integrates aspects of overlap, distance, and aspect ratio consistency to measure the difference between predicted and ground truth bounding boxes. As a result, it enables the model to more precisely target the object’s shape, size, and orientation.
For the segmentation tasks, we utilize an identical loss function. That means $\mathcal{L}_{\text{segda}}$ has the same formula with $\mathcal{L}_{\text{segll}}$. We will collectively refer to them as $\mathcal{L}_{\text{seg}}$. The formula is shown as follows:
@@
where $\mathcal{L}_{\text{FL}}$ and $\mathcal{L}_{\text{TL}}$ are focal loss [24] and tversky loss [32]. Both are widely used loss functions in segmentation tasks. $\lambda_{\text{FL}}$ and $\lambda_{\text{TL}}$ are corresponding coefficients.
$\mathcal{L}_{\text{FL}}=-\alpha_{t}\left(1-p_{t}\right)^{\gamma}\log\left(p_{t}\right)$ (7)
where $p_{t}$ is the probability of the model predicting the positive class. $y$ is the ground truth of each pixel. $\alpha_{t}$ is a weighting factor to balance the importance of positive/negative examples. $\gamma$ is a focusing parameter to modulate the contribution of each example to the loss. Focal loss offers a robust solution for handling imbalance samples, ensuring that the model doesn’t become overwhelmingly biased towards the predominant and easily learned class. Instead, it places greater emphasis on challenging and underrepresented areas.
$\mathcal{L}_{\text{TL}}=1-\frac{TP}{TP+\alpha FN+\beta FP}$ (8)
Tversky loss is an extension of the Dice loss, introducing two additional parameters ($\alpha$ and $\beta$) to assign distinct weights to false positives and false negatives, thereby enhancing its capability to handle imbalance tasks.
@@ III-D Training Paradigm
Our training paradigm is different from the multi-task learning methods prevalent in panoptic autonomous driving tasks. We utilize an end-to-end training mode, performing backward only once for each batch. This means the entire network is optimized collectively, without freezing specific layers or alternating optimization, thus reducing the training time. Algorithm 3 illustrates the step-by-step training process. In each epoch, the required predicted results list $\hat{y}$ is obtained through a single forward propagation, encompassing information such as detection bounding boxes, classification information, and segmentation masks. Subsequently, each task loss is calculated and summed into a single loss $\mathcal{L}$. Then, backward propagation is performed only once to optimize the model across all tasks. After completing an epoch of training, the model is evaluated. If there’s no improvement in performance compared to the last 50 epochs, training terminates prematurely. Otherwise, the training will stop after training 300 epochs.
@@
0:  Target end-to-end network $F$ with parameters $\Theta$; Training dataset $\tau$; Validation dataset $x_{val}$; Threshold for convergence $\phi$; $y$ is the label list
@@
4:        $\hat{y}\leftarrow F(x)$ $\hat{y}$ is a prediction list
5:        $\mathcal{L}_{\text{det}}\leftarrow(y[0],\hat{y}[0])$
6:        $\mathcal{L}_{\text{segda}}\leftarrow(y[1],\hat{y}[1])$
7:        $\mathcal{L}_{\text{segll}}\leftarrow(y[2],\hat{y}[2])$
8:        $\mathcal{L}=\mathcal{L}_{\text{det}}+\mathcal{L}_{\text{segda}}+\mathcal{L}_{\text{segll}}$
9:        $\Theta\leftarrow\arg\min_{\Theta}\mathcal{L}$
10:     end for
11:     $p\leftarrow F(x_{val})$
12:     if $p=p_{\text{epoch}-50}$ then
13:        return  $F(x;\Theta)$ If the model hasn’t any improvements over the last 50 epochs, then training terminates prematurely.
14:     end if
15:  end for
16:  return  $F(x;\Theta)$
@@ IV EXPERIMENTS AND RESULTS
In this section, we evaluate our model’s performance and inference time on the BDD100K dataset and compare it against classical methods used in multi-task autonomous driving panoptic perception. Additionally, we present extensive ablation studies and provide an analysis of the experimental results.
@@ IV-A1 Dataset
The BDD100K dataset is a prominent resource in the study of autonomous driving, including 100k samples and multi-task annotations. Beyond sheer volume, the dataset’s significance is underscored by its multifaceted nature, comprising diverse geographies, environmental contexts, and weather conditions. These advantages ensure that models trained on BDD100K achieve robustness and versatility, making it an ideal choice for our research. The dataset is divided into three parts: a training set with 70k images, a validation set of 10k images, and a testing set encompassing 20k images. Since the labels for the testing set are not public, we evaluate our model on the validation set. Similar to YOLOP, our detection task focuses solely on “vehicle” detection, encompassing categories including car, bus, truck, and train.
@@ IV-A2 Evaluation Metrics
We utilize recall and mAP50 as the evaluation metrics for object detection tasks. Both metrics are widely recognized and accepted in the detection task. Recall indicates a model’s capability to detect all object instances from the designated classes accurately. mAP50 is derived by taking the mean Average Precision for all classes at an IoU threshold of 0.5. Notably, Average Precision (AP) quantifies the area beneath the precision-recall curve. For the segmentation tasks, similar to YOLOP [16], we utilize mIoU to evaluate the drivable area segmentation task. For the lane line segmentation task, we employ both accuracy and IoU as evaluation metrics. However, due to the number of pixel imbalances between the background and foreground in lane line segmentation, we adopt a more meaningful balance accuracy metric for evaluation. Traditional accuracy will skew results by favouring classes with a larger number of samples. In contrast, balance accuracy offers a more equitable metric by considering each class’s accuracy. The formula is as follows:
Line $\text{Line Accuracy}=\frac{\text{Sensitivity}+\text{Specificity}}{2}$ (9)
@@ IV-A3 Experimental Setup and Implementation
We compare our model with several prominent multi-task models. Due to the scarcity of multi-task models, we also include outstanding single-task or two-task models in our comparison. YOLOP and HybridNet represent the state-of-the-art and are open-source multi-task models in panoptic autonomous driving, specifically on the BDD100K dataset. Faster R-CNN and YOLOv8 are exemplary representations of two-stage and one-stage object detection networks, respectively. Both MultiNet and DLT-Net could address multiple panoptic driving perception tasks, and they have demonstrated good performance in object detection and drivable area segmentation on the BDD100k dataset. PSPNet [33] excels in semantic segmentation tasks due to its unparalleled capability to aggregate global information. Given the absence of an appropriate multi-task network handling lane line segmentation on the BDD100K dataset, we compare our model against Enet [34], SCNN, and Enet-SAD [35], which are three leading lane detection networks.
@@
We train our model using the SGD optimizer with a learning rate ($lr$) of 0.01, momentum of 0.937, and a weight decay of 0.0005. Initially, our model undergoes a warm-up training for 3 epochs. During this warm-up phase, the momentum of the SGD optimizer is set to 0.8, and the bias lr is 0.1. In our training process, we adopt a linear learning rate annealing strategy. This strategy helps ensure that the model learns rapidly in the early stages of training and converges more stably in the later stages. Additionally, we resize the original image dimensions from 1280 × 720 to 640 × 640. For loss function coefficients, we set $\lambda_{\text{FL}}=24.0$, $\lambda_{\text{TL}}=8.0$, $\lambda_{\text{DFL}}=1.5$, $\lambda_{\text{CIoU}}=7.5$, and $\lambda_{\text{BCE}}=0.5$. We adopt $\alpha=0.7$ and $\beta=0.3$ in $\mathcal{L}_{\text{TL}}$, $\alpha_{t}=0.25$ and $\gamma=2$ in $\mathcal{L}_{\text{FL}}$. Finally, we train with a batch size of 120 on three RTX 4090 for 300 epochs. For evaluation, we set the confidence threshold at 0.001 and the Non-Maximum Suppression (NMS) threshold at 0.6. For prediction, the confidence threshold is set at 0.25, and the NMS threshold is at 0.45. We follow the confidence threshold and NMS settings from YOLOP. As a result, the visualization may slightly differ from the quantitative results.
@@ IV-B Experimental results
In this subsection, we train our model end-to-end and compare its performance to other outstanding methods.
@@ IV-B1 Inference Time
One of the primary challenges in deep learning applications is the inference time. Especially in autonomous driving tasks, require deploying models on edge devices, which usually have limited computational resources. Thus, ensuring models are both lightweight and real-time becomes paramount. In Table I, we reproduce and test the FPS of YOLOP [^1] https://github.com/hustvl/YOLOP, HybridNet [^2] https://github.com/datvuthanh/HybridNets, YOLOv8333https://github.com/ultralytics/ultralytics, and our model, all FPS tests on a GTX 1080 Ti GPU with 1 and 32 batch sizes. The FPS calculation approach follows HybridNet. Additionally, we provide the number of parameters for each model as one of the evaluation results.
@@
Model Parameters FPS (bs=1) FPS (bs=32)
YOLOP 7.9M 26.0 134.8
HybridNet 12.83M 11.7 26.9
YOLOv8n(det) 3.16M 102 802.9
YOLOv8n(seg) 3.26M 82.55 610.49
A-YOLOM(n) 4.43M 39.9 172.2
A-YOLOM(s) 13.61M 39.7 96.2
The primary difference between A-YOLOM(n) and A-YOLOM(s) is the complexity of the backbone. A-YOLOM(n) is our lightweight backbone network designed with reduced complexity, making it ideal for deployment on edge devices. A-YOLOM(s) has a more complex backbone, offering more powerful performance with increased time overhead, especially in scenarios with bs=32. Compared to other SOTA multi-task models and single-task models in panoptic autonomous driving, A-YOLOM(n) stands out as more lightweight and offers more efficiency. Specifically, compared to YOLOP, A-YOLOM(n) has significantly fewer parameters and a higher FPS. It achieves a speed-up of x1.53 at bs=1 and x1.28 at bs=32. That indicates our model is more efficient. Compared to HybridNet, A-YOLOM(n) has a significant advantage in parameters and speed. Despite having more parameters than HybridNet, A-YOLOM(s) is faster. HybridNet doesn’t meet the requirements for real-time performance in both bs=1 and bs=32. We believe the longer inference time for HybridNet is due to its being an anchor-based method. Generating numerous anchor boxes will increase computational overhead, consequently reducing the overall speed of inference. In this case, we won’t compare its performance with ours in the subsequent section. YOLOv8 is a single-task model, meaning it can only implement one task within a single model. This is the reason we have listed two versions: YOLOv8n(det) and YOLOv8n(seg). Although both of them are faster than other models, including ours, they require deploying three separate models on edge devices, which means combining one detection model with two segmentation models into one edge device, amounting to 9.68M parameters. This is x2.18 more than A-YOLOM(n), putting significant pressure on edge devices. Additionally, YOLOv8’s performance across the three tasks is much inferior to ours, which we will discuss in the upcoming section.